On the way to school, the bus speeds up from 20 m/s to 36 m/s in 4 seconds. What distance does the bus cover in this time frame?
1
Expert's answer
2019-07-09T13:12:20-0400

We can find the distance covered by the bus from the kinematic equation:


"d = \dfrac{v_i + v_f}{2}t,"

here, "d" is the distance covered by the bus, "v_i = 20 m\/s" is the initial velocity of the bus, "v_f = 36 m\/s" is the final velocity of the bus, "t = 4 s" is time during which the bus speeds up.

Then, from this formula we can calculate the distance covered by the bus:


"d = \dfrac{v_i + v_f}{2}t = \dfrac{20 \dfrac{m}{s} + 36 \dfrac{m}{s}}{2} \cdot 4 s = 112 m."

Answer:

"d = 112 m."